OpenStudy (anonymous):

The line passes through the points ( 2, 1) and ( 5, 4)

OpenStudy (perl):

thats a statement, not a question :)

OpenStudy (anonymous):

The line passes through the points ( 2, 1) and ( 5, 4) a. y = x + 1 c. y = x - 1 b. y = -x - 1 d. y = x + 3

OpenStudy (midhun.madhu1987):

 this is the equation of a line passing through 2 points...